Suppose for a moment there were a million cars, only one of them green, the rest blue. If the witness said the car was green, would you think it more likely that the car was green, or that the witness made a mistake?
Suppose further that there was one green car out of a million, and that the witness was actually just guessing, and had a 50% chance to say green or blue. This means the actual colour of the car is irrelevant to the witness's testimony.
Before the witness said anything about the colour of the car, you would say there's a 50% chance the witness will get it right. But if the witness said the car was green, would you say the same thing?
The witness is still (by definition) correct 80% of the time - it will never change by definition. Your ability to predict the which selection is correct does not affect his ability to be correct 80% of the time. You can predict in this exclusive sample set that the remaining guesses will be blue if the reliability is accurate but the actual selection doesn't change his reliability. As long as there is a viable sample to choose from independent of the ratio of individual objects - there is an 80% chance he/she will be correct. You can use the numbers in the pool to predict results but his 80% correctness will never vary.
I think
What we're talking about here is called a conditional probability. In simple terms, it means that if you know something about the outcome of some random experiment, and you may be able to "update" your understanding of the probability of certain outcomes
under these conditions.
I'm not saying the reliability of the witness changes in a general sense. I'm saying that if you know what the witness said about the colour of the car, then you can use that information (along with the distribution of colours among the cars) to know the probability the witness is correct in this situation.
It might be easier to see an example of conditional probability in another situation. Suppose I roll two dice and add them together, but hide the roll from you. What is the probability that the total is 12, based on what you know? 1/36, because you need to roll 6 on both dice, each of which has a probability of 1/6, and 1/6 * 1/6 = 1/36. But now suppose I reveal the first rolled die to you. If that die is a 6, you can now say the probability that the total is 12 is 1/6, because you now only need the second die to be a 6. On the other hand, if the first die is not a 6, then the probability that the total is 12 is 0. It isn't possible. These are two examples of conditional probability, the probability that the total is 12
given that the first die is 6 or the probability that the total is 12
given that the first die is not 6. You aren't saying the probability of a particular roll has changed in general, but that the probabilities are different when you already know something about what happened.
This stuff is involved in what I do for a living, by the way. Not that that necessarily means anything; I've certainly met people that do something involving this stuff for a living that don't actually understand it.